# Decoupled Multimodal Fusion for User Interest Modeling in Click-Through Rate Prediction

### v1

@@ 4. Methodology
The framework of DMF is illustrated in Fig. 1. Due to the semantic gap between ID and multimodal embeddings, we use target-aware similarity scores as alignment signals. These features are fed into both the Similarity Histogram and DTA modules to derive complementary interest representations. In particular, the constructed features are transformed into learnable embeddings by the MSE module to enable computational decoupling within the DTA module. The CMM module fuses these representations to capture both semantic generalization and fine-grained behavioral patterns. The final representation is combined with auxiliary features (e.g., user profile embeddings) and fed into an MLP for prediction.
4.1. Target-aware Multimodal Similarity Feature
@@
Images and text descriptions capture rich semantic signals that reflect item relationships, which we leverage as multimodal inputs. We use RoBERTa and Vision Transformer (ViT) to encode item titles and main images into textual and visual representations, respectively. These representations are fused and used to fine-tune the encoders via contrastive learning on item pairs, encouraging semantically similar items to have aligned embeddings. The resulting frozen embeddings are used to compute multimodal similarity scores in downstream modeling.
@@
As mentioned before, direct feature fusion with raw embeddings of multimodal information is ineffective for CTR prediction due to distinct spaces. State-of-the-art multimodal recommenders often freeze pretrained modality encoders and leverage item similarity computed by multimodal representations as auxiliary input to guide user interest modeling. Given a user $u$, the historical interaction sequence is $\mathcal{X}_{u}=[N_{1},N_{2},\ldots,N_{L}]$, where $L$ is the sequence length, $N_{c}$ is the target item. The similarity between the target item $N_{c}$ and each historical item $N_{i}$ is computed via cosine similarity on their multimodal embeddings:
(5) $S_{i}=\text{sim}(v_{c},v_{i})=\frac{v_{c}^{\top}v_{i}}{\|v_{c}\|\|v_{i}\|},\quad\forall i\in\{1,\dots,L\},$
where $v_{c},v_{i}\in\mathbb{R}^{d}$ denote the frozen multimodal representations of $N_{c}$ and $N_{i}$, respectively.
The multimodal similarity score is computed based on the content alignment between the candidate item and each historical interaction. Due to its dependency on the candidate item, this multimodal similarity score could be characterized as a target-aware feature.
4.2. Target-Aware and Target-Agnostic Nodes
@@
In the computation graph of a CTR model, we identify two types of intermediate nodes based on their input features and tensor shapes: target-aware nodes and target-agnostic nodes.
@@
A target-aware node processes target-aware features for a batch of candidate items, and thus has a leading dimension of $B$, where $B$ denotes the number of items being scored. In contrast, a target-agnostic node processes user-level features (e.g., demographics, historical interactions) and outputs a $1\times d$ tensor computed once per user. For example, in a 2-dimensional tensor representation, a target-agnostic node outputs a tensor of size $1\times d$, while an item-side node outputs $B\times d$. Due to the shape difference, target-agnostic nodes are often more computationally efficient, as the computational overhead does not scale with the number of candidates [^1] When a downstream operator (e.g., concatenation or element-wise addition) takes inputs from both a target-agnostic node and a target-aware node, broadcast-compatible operators can be used to align the shapes of these two nodes, avoiding redundant computation caused by explicit replication.. The reusability property of target-agnostic nodes has been widely exploited in industrial systems to optimize online inference complexity (Zhai et al., 2024).
4.3. Decoupled Target Attention
Compared with DIN, the online inference complexity of MHTA can be greatly optimized by applying the reusability property of target-agnostic nodes to the linear projections of K and V. DIN cannot benefit from this optimization due to the early fusion between the target items and sequential interactions before being fed into the MLP. Therefore, we employ MHTA as the target-aware attention approach and integrate the target-aware multimodal similarity feature constructed in Section 4.1 into MHTA.
A straightforward way for integrating target-aware multimodal similarity features $S_{sim}=[S_{1},S_{2},\ldots,S_{L}]$ is to treat the features as side information and feed them together with ID-based embeddings into MHTA. As shown in Fig. 2(a), we denote this method as an early fusion method. By incorporating ID-based embeddings and similarity scores at the input level, the key (K) and value (V) become target-aware nodes. Prior work shows that target-aware computation tends to yield better model performance when integrated earlier into the sequential model structure (Xia et al., 2023). However, the prohibitive computational overhead is unacceptable for industrial recommendation systems with more than $10^{3}$ candidate items. In contrast, later fusion delays the integration of ID-based embeddings and similarity scores until the model’s final layer as in Fig. 2(b). The K and V computed by ID-based embeddings become target-agnostic nodes, and the calculations can be reused. As the late fusion method utilizes a modality-centric modeling strategy, it fails to capture fine-grained interactions between content semantics and behavioral signals. All these weaknesses motivate us to find a fusion model for target-aware features that balances efficiency and effectiveness.
In this paper, we propose Decoupled Target Attention (DTA), which decouples the computation of the K based on ID-derived embeddings from that based on multimodal similarity features, as shown in Fig. 2(c). The same decoupling is applied to the computation of the V. For the ID part, the K and V become target-agnostic nodes, and the calculations can be reused. For the target-aware features part, we design a Multimodal Similarity Encoding (MSE) module to transform similarity scores into learnable embeddings. The module processes raw cosine similarities by discretizing them into buckets and performing embedding lookup, enabling differentiable and efficient integration of target-aware side information. This avoids the need for linear projection layers when computing key and value representations (denoted as $K^{\prime}$ and $V^{\prime}$ in Fig. 2(c)), which would otherwise require $B$-fold recomputation during online inference. After that, $K$ and $K^{\prime}$ are fused via element-wise addition, and the same operation is applied to $V$ and $V^{\prime}$, enabling fine-grained interactions between ID-based collaborative representations and multimodal representations.
To formalize this design, let $N_{id}\in\mathbb{R}^{L\times d}$ denote the ID-based embeddings of the user’s historical interaction sequence, and $S=[S_{1},\dots,S_{L}]\in\mathbb{R}^{L}$ the multimodal similarity scores between the target item and each historical item. Our goal is to compute query $Q^{D}$, key $K^{D}$, and value $V^{D}$ for target attention, where $K^{D}$ and $V^{D}$ are constructed by combining ID-derived components (target-agnostic) and similarity-enhanced components (target-aware).
Specifically, MSE performs the operations of discretization and embedding lookup to obtain the similarity embeddings. The range of the multimodal similarity scores is [1.0,1.0][-1.0,1.0]. We partition the score range into several predefined buckets based on their absolute values via equal frequency discretization. By binning the multimodal similarity scores $S$ as codes and mapping codes to $D$-dimensional vectors, we obtain $S^{D}=[S_{1}^{D},S_{2}^{D},S_{3}^{D},\ldots,S_{T}^{D}]$. Instead of projecting similarity scores through linear layers (which would require B-times computation), we discretize scores into buckets and perform embedding lookup, a constant-time operation independent of B. Subsequently, we add $S_{D}$ to $K_{D}$ and $V_{D}$ to form the inputs for the scaled dot-product attention. The output vector from the cross-features fusion target attention is then concatenated with other embeddings and fed into the following decision layers for CTR prediction:
(6) $Q^{D}=\text{W}_{Q}(I),$
(7) $S_{K}^{D},S_{V}^{D}=Lookup_{K}(Bucket(S)),Lookup_{V}(Bucket(S)),$
(8) $K^{D}=\text{W}_{K}^{D_{nid}}(N_{id})+S_{K}^{D},$
(9) $V^{D}=\text{W}_{V}^{D_{nid}}(N_{id})+S_{V}^{D},$
(10) $Output=TargetAttention(Q^{D},K^{D},V^{D}),$
where $N_{id}$ is the ID-based embeddings of user interaction sequences.
4.4. Complementary Modality Modeling
@@
Figure 2. Comparison of side information fusion methods based on target-aware attention: (a) early fusion, (b) late fusion, and (c) decoupled fusion, which balances efficiency and effectiveness.:
@@
The modality-centric and modality-enriched modeling strategies capture user interests from fundamentally different perspectives, making them complementary. The modality-centric modeling strategy answers whether an item is semantically consistent with the user’s past interests, supporting broad generalization, while the modality-enriched modeling strategy determines which among the semantically related items has been most behaviorally significant, enabling precise personalization. Their integration allows for benefiting from both stable semantic matching and fine-grained interaction modeling.
@@
For modality-centric modeling, we employ a histogram-based approach (Sheng et al., 2024) to obtain user interest representations. Given the multimodal similarity scores, the score range from -1.0 to 1.0 is uniformly partitioned into N predefined intervals. For each interval, the number of similarity scores falling within that range is counted. As a result, an N-dimensional histogram vector is constructed, where each dimension corresponds to the frequency of scores in the respective bin. Then, the N-dimensional histogram vector is fed into a subsequent multilayer perceptron for further processing. For modality-enriched modeling, we exploit DTA to obtain user interest representations, detailed in Section 4.3.
@@
The final representation $\text{R}_{u}$ is obtained by aggregating the user interest representations from modality-centric and modality-enriched modeling, respectively, as shown in Eq. (11). While modality-centric modeling strategies emphasize semantic coherence through content-driven generalization, independent of item identities, modality-enriched modeling strategies capture fine-grained behavioral preferences by integrating identity signals with multimodal context, thereby prioritizing context-aware personalization.
(11) $\text{R}_{u}=\alpha\text{R}_{me}+(1-\alpha)\text{R}_{mc},$

### v2

@@ IV Methodology
The framework of DMF is illustrated in Fig. 2. Due to the semantic gap between ID-based embeddings and multimodal representations, we use target-aware similarity scores as alignment signals. These features are fed into both the Similarity Histogram and DTA modules to derive complementary interest representations. In particular, the constructed features are transformed into learnable embeddings to enable computational decoupling within the DTA module. The CMM module fuses these representations to capture both semantic generalization and fine-grained behavioral patterns. The final representation is combined with auxiliary features (e.g., user profile embeddings) and fed into an MLP for prediction.
IV-A Target-aware Multimodal Similarity Feature
@@
Images and text descriptions capture rich semantic signals that reflect item relationships, which we leverage as multimodal inputs. We use RoBERTa [34] and Vision Transformer (ViT) [35] to encode item titles and main images into textual and visual representations, respectively. These representations are fused and used to fine-tune the encoders via contrastive learning on item pairs, encouraging semantically similar items to have aligned embeddings. The resulting frozen embeddings are used to compute multimodal similarity scores in downstream modeling.
@@
As previously discussed, direct feature fusion with raw embeddings of multimodal information is ineffective for CTR prediction due to distinct spaces. State-of-the-art multimodal recommenders often freeze pretrained modality encoders and leverage item similarity computed by multimodal representations as auxiliary input to guide user interest modeling. Let $\mathbf{m}_{c}\in\mathbb{R}^{d_{m}}$ and $\mathbf{m}_{j}\in\mathbb{R}^{d_{m}}$ denote the frozen multimodal representations of the target item and the $j$-th historically interacted item, respectively, where $d_{m}$ is the dimension of the multimodal representation. The multimodal similarity score between them is computed using cosine similarity:
$c_{j}=\frac{\mathbf{m}_{c}^{\top}\mathbf{m}_{j}}{\|\mathbf{m}_{c}\|_{2}\|\mathbf{m}_{j}\|_{2}},\quad\forall j\in\{1,\dots,L\}.$ (10)
We define the similarity vector as $\mathbf{c}_{\text{sim}}=[c_{1},c_{2},\ldots,c_{L}]^{\top}\in\mathbb{R}^{L}$, where each $c_{j}\in[-1,1]$ denotes the cosine similarity between the target item and the $j$-th historically interacted item. This vector serves as target-aware side information in subsequent modeling.
IV-B Target-Aware and Target-Agnostic Nodes
@@
In the computation graph of a CTR model, we classify intermediate nodes into two categories based on their input features and tensor shapes, i.e., target-aware nodes and target-agnostic nodes. As illustrated in Fig. 3, this distinction has direct implications for online inference efficiency.
Figure 3: Illustration of target-agnostic and target-aware node computation in a CTR model. Left panel shows target-agnostic features derived from user interaction sequence, which are computed once and reused across all target items. Right panel shows target-aware features, where each target item triggers a distinct computation path for its associated historical interactions. The output of target-agnostic nodes is reusable, while target-aware nodes require recomputation per target item
@@
A target-aware node processes target-aware features for a batch of target items, and thus has a leading dimension of $B$, where $B$ denotes the number of items being scored. In contrast, a target-agnostic node processes user-level features (e.g., demographics, historically interacted items) and outputs a $1\times d$ tensor computed once per user. For example, in a 2-dimensional tensor representation, a target-agnostic node outputs a tensor of size $1\times d$, while an item-side node outputs $B\times d$. Due to the shape difference, target-agnostic nodes are often more computationally efficient, as the computational overhead does not scale with the number of target items∗*∗**When a downstream operator (e.g., concatenation or element-wise addition) takes inputs from both a target-agnostic node and a target-aware node, broadcast-compatible operators can be used to align the shapes of these two nodes, avoiding redundant computation caused by explicit replication.. The reusability property of target-agnostic nodes has been widely exploited in industrial systems to optimize online inference complexity [36].
IV-C Decoupled Target Attention
@@
Figure 4: Comparison of side information fusion methods based on target-aware attention: (a) early fusion, (b) late fusion, and (c) decoupled fusion, which balances efficiency and effectiveness.
Compared with DIN, the online inference complexity of MHTA can be greatly optimized by applying the reusability property of target-agnostic nodes to the linear projections of $\mathbf{K}$ and $\mathbf{V}$. DIN cannot benefit from this optimization due to the early fusion between the target items and sequential interactions before being fed into the MLP. Therefore, we employ MHTA as the target-aware attention approach and integrate the target-aware multimodal similarity feature constructed in Section IV-A into MHTA.
A straightforward way for integrating the target-aware multimodal similarity vector $\mathbf{c}_{\text{sim}}=[c_{1},c_{2},\ldots,c_{L}]^{\top}$ is to treat it as side information and feed it together with ID-based embeddings into MHTA. As shown in Fig. 4(a), this approach is referred to as early fusion. By incorporating ID-based embeddings and similarity scores at the input level, the key matrix $\mathbf{K}$ and value matrix $\mathbf{V}$ become target-aware nodes. Prior work shows that target-aware computation tends to yield better model performance when integrated earlier into the sequential model structure [6]. However, the prohibitive computational overhead is unacceptable for industrial recommendation systems with more than $10^{3}$ target items. Late fusion defers the integration of ID-based embeddings and similarity scores until the final prediction layer, as illustrated in Fig. 4(b). In late fusion, ID-based embeddings and multimodal similarity scores are processed separately, where the ID-based embeddings are projected via linear layers to obtain $\mathbf{K}_{\mathrm{id}}$ and $\mathbf{V}_{\mathrm{id}}$, while the similarity scores $\{c_{j}\}$ are similarly projected to produce auxiliary key and value matrices $\mathbf{K}_{\mathrm{sim}}$ and $\mathbf{V}_{\mathrm{sim}}$. The $\mathbf{K}_{\mathrm{id}}$ and $\mathbf{V}_{\mathrm{id}}$ computed by ID-based embeddings become target-agnostic nodes, and the calculations can be reused. As the late fusion method utilizes a modality-centric modeling strategy, it fails to capture fine-grained interactions between content semantics and behavioral signals. All these weaknesses motivate us to find a fusion model for target-aware features that balances efficiency and effectiveness.
In this paper, we propose Decoupled Target Attention (DTA), which decouples the computation of the $\mathbf{K}$ based on ID-derived embeddings from that based on multimodal similarity features, as shown in Fig. 4(c). DTA extends MHTA by augmenting its key and value projections with modality-aware signals, while preserving the target-agnostic structure of MHTA for $\mathbf{K}_{\mathrm{id}}$ and $\mathbf{V}_{\mathrm{id}}$. The same decoupling applies to the computation of the value matrix $\mathbf{V}$. For the ID-derived components, the key matrix $\mathbf{K}_{\mathrm{id}}$ and value matrix $\mathbf{V}_{\mathrm{id}}$ are target-agnostic and thus reusable across different target items. For the multimodal similarity features, each scalar similarity score $c_{j}$ is discretized into a bucket index $b_{j}=\mathrm{bucket}(c_{j})$, and then mapped into separate learnable embedding vectors for keys and values, where $\mathbf{e}_{b_{j}}^{K},\mathbf{e}_{b_{j}}^{V}\in\mathbb{R}^{d_{h}}$. The raw cosine similarities are processed by discretizing them into buckets and performing embedding lookup, enabling differentiable and efficient integration of target-aware side information. This avoids linear projection layers when computing auxiliary key $\mathbf{K}_{\mathrm{sim}}$ and value $\mathbf{V}_{\mathrm{sim}}$, which would otherwise necessitate $B$ times recomputation per user sequence during online inference.
To formalize this design, let $\mathbf{S}=[\mathbf{s}_{1},\mathbf{s}_{2},\ldots,\mathbf{s}_{L}]^{\top}\in\mathbb{R}^{L\times d}$ denote the ID-based embedding matrix of the historical interaction sequence, and let $\mathbf{c}_{\text{sim}}=[c_{1},c_{2},\ldots,c_{L}]^{\top}\in\mathbb{R}^{L}$ denote the multimodal similarity scores between the target item and each historical item, as defined in Eq. (10). Our goal is to compute the query $\mathbf{Q}^{\mathrm{DTA}}\in\mathbb{R}^{B\times d_{h}}$, key $\mathbf{K}^{\mathrm{DTA}}\in\mathbb{R}^{L\times d_{h}}$, and value $\mathbf{V}^{\mathrm{DTA}}\in\mathbb{R}^{L\times d_{h}}$ for DTA.
Specifically, each similarity score $c_{j}\in[-1,1]$ is discretized into one of $M$ predefined buckets using equal frequency discretization [37], producing a bucket index $b_{j}\in\{1,\dots,M\}$. Let $b_{j}=\text{bucket}(c_{j})$ denote the bucket index of $c_{j}$. Each bucket $m\in\{1,\dots,M\}$ corresponds to two independent learnable embedding vectors $\mathbf{e}_{m}^{K},\mathbf{e}_{m}^{V}\in\mathbb{R}^{d_{h}}$, forming two lookup tables $\{\mathbf{e}_{m}^{K}\}_{m=1}^{M}$ and $\{\mathbf{e}_{m}^{V}\}_{m=1}^{M}$. We define the similarity embedding matrices as:
$\mathbf{E}_{\mathrm{sim}}^{K}$ $=[\mathbf{e}_{b_{1}}^{K},\mathbf{e}_{b_{2}}^{K},\ldots,\mathbf{e}_{b_{L}}^{K}]^{\top}\in\mathbb{R}^{L\times d_{h}},$ (11)
$\mathbf{E}_{\mathrm{sim}}^{V}$ $=[\mathbf{e}_{b_{1}}^{V},\mathbf{e}_{b_{2}}^{V},\ldots,\mathbf{e}_{b_{L}}^{V}]^{\top}\in\mathbb{R}^{L\times d_{h}}.$ (12)
Instead of projecting similarity scores through linear layers (which would require $B$-times computation), we discretize scores into buckets and perform embedding lookup, a constant-time operation independent of $B$. Subsequently, $\mathbf{E}_{\mathrm{sim}}^{K}$ and $\mathbf{E}_{\mathrm{sim}}^{V}$ are added to $\mathbf{K}_{\mathrm{id}}$ and $\mathbf{V}_{\mathrm{id}}$, respectively, to form the final key and value matrices for scaled dot-product attention.
To ensure that the decoupled fusion strategy does not sacrifice expressive power compared to the early fusion approach, we formally establish a condition under which both methods are computationally equivalent. Specifically, we show that when the discretization scheme employs a coarse bucketing strategy with sufficient capacity, the output of decoupled fusion can approximate that of early fusion to arbitrary precision.
Theorem 1 (Approximate Expressiveness of Decoupled Fusion). To illustrate the expressiveness of our discretization scheme, we consider a simplified early fusion model where the hidden representation of the $i$-th historical item is computed as:
$\mathbf{h}_{i}^{\prime}=\mathbf{s}_{i}+c_{i}\cdot\mathbf{w},$ (13)
where $\mathbf{s}_{i}\in\mathbb{R}^{d}$ is the ID-based embedding of the $i$-th historically interacted item, $c_{i}\in[-1,1]$ denotes its multimodal similarity score with respect to the target item, and $\mathbf{w}\in\mathbb{R}^{d}$ is a shared learnable parameter vector that modulates the influence of the similarity signal.
In contrast, the proposed decoupled fusion computes:
$\mathbf{h}_{i}^{\text{final}}=\mathbf{s}_{i}+\mathbf{e}_{b_{i}},$ (14)
where $b_{i}=\text{Bucket}(c_{i})$ maps $c_{i}$ to one of $M$ discrete buckets, and $\mathbf{e}_{b_{i}}\in\mathbb{R}^{d_{h}}$ is a learnable embedding associated with bucket $b_{i}$.
Then, for any arbitrarily small approximation error $\epsilon>0$, there exists a sufficiently large number of buckets $M$ and an appropriate choice of $\{\mathbf{e}_{m}\}_{m=1}^{M}$ such that:
$\max_{i}\|\mathbf{h}_{i}^{\prime}-\mathbf{h}_{i}^{\text{final}}\|_{2}<\epsilon.$ (15)
Theorem 1 justifies our design choice. Through a coarse-bucketing scheme endowed with sufficient capacity, DTA attains nearly lossless expressiveness while delivering substantial computational efficiency. Despite its simplicity, the discretization with learnable embeddings effectively captures complex patterns in the similarity signal, preserving the modeling power of decoupled fusion.
Let $\mathbf{S}=[\mathbf{s}_{1},\mathbf{s}_{2},\ldots,\mathbf{s}_{L}]^{\top}\in\mathbb{R}^{L\times d}$ denote the ID-based embedding matrix of the historical interaction sequence. We compute the query, key, and value for Decoupled Target Attention (DTA) as follows. The query $\mathbf{Q}^{\mathrm{DTA}}\in\mathbb{R}^{B\times d_{h}}$ is derived from the target item embeddings $\mathbf{q}_{i}$ for each target item:
$\mathbf{q}_{i}^{\mathrm{DTA}}=\mathbf{W}_{q}\mathbf{q}_{i},\quad\mathbf{Q}^{\mathrm{DTA}}=[\mathbf{q}_{1}^{\mathrm{DTA}},\ldots,\mathbf{q}_{B}^{\mathrm{DTA}}]\in\mathbb{R}^{B\times d_{h}},$ (16)
where $\mathbf{W}_{q}\in\mathbb{R}^{d_{h}\times d}$ is a learnable projection matrix.
The base key and value matrices are computed from ID-based embeddings:
$\mathbf{K}_{\mathrm{id}}$ $=\mathbf{S}\mathbf{W}_{k}^{\top}\in\mathbb{R}^{L\times d_{h}},$ (17)
$\mathbf{V}_{\mathrm{id}}$ $=\mathbf{S}\mathbf{W}_{v}^{\top}\in\mathbb{R}^{L\times d_{h}},$ (18)
where $\mathbf{W}_{k},\mathbf{W}_{v}\in\mathbb{R}^{d_{h}\times d}$ are learnable projection matrices. These matrices are independent of the target item and thus reusable across multiple target items.
Each similarity score $c_{j}$ is mapped to separate learnable embeddings $\mathbf{e}_{b_{j}}^{K}$ and $\mathbf{e}_{b_{j}}^{V}$, producing the similarity embedding matrices:
$\mathbf{E}_{\mathrm{sim}}^{K}$ $=[\mathbf{e}_{b_{1}}^{K},\mathbf{e}_{b_{2}}^{K},\ldots,\mathbf{e}_{b_{L}}^{K}]^{\top}\in\mathbb{R}^{L\times d_{h}},$ (19)
$\mathbf{E}_{\mathrm{sim}}^{V}$ $=[\mathbf{e}_{b_{1}}^{V},\mathbf{e}_{b_{2}}^{V},\ldots,\mathbf{e}_{b_{L}}^{V}]^{\top}\in\mathbb{R}^{L\times d_{h}},$ (20)
with dimensionality $d_{h}$ matching the attention hidden size.
The final key and value matrices are obtained by element-wise addition:
$\mathbf{K}^{\mathrm{DTA}}$ $=\mathbf{K}_{\mathrm{id}}+\mathbf{E}_{\mathrm{sim}}^{K},$ (21)
$\mathbf{V}^{\mathrm{DTA}}$ $=\mathbf{V}_{\mathrm{id}}+\mathbf{E}_{\mathrm{sim}}^{V},$ (22)
ensuring fine-grained interaction while preserving reusability of $\mathbf{K}_{\mathrm{id}}$ and $\mathbf{V}_{\mathrm{id}}$.
The output of the DTA module is computed as:
$\mathbf{U}^{\mathrm{DTA}}=\mathrm{DTA}(\mathbf{Q}^{\mathrm{DTA}},\mathbf{K}^{\mathrm{DTA}},\mathbf{V}^{\mathrm{DTA}})\in\mathbb{R}^{B\times d_{h}},$ (23)
where $\mathbf{U}^{\mathrm{DTA}}\in\mathbb{R}^{B\times d_{h}}$ is the resulting user interest representation.
IV-D Complementary Modality Modeling
@@
The modality-centric and modality-enriched modeling strategies capture user interests from fundamentally different perspectives, making them complementary. The modality-centric modeling strategy answers whether an item is semantically consistent with user interests, supporting broad generalization, while the modality-enriched modeling strategy determines which among the semantically related items has been most behaviorally significant, enabling precise personalization. Their integration allows for benefiting from both stable semantic matching and fine-grained interaction modeling.
@@
For modality-centric modeling, we employ a similarity histogram-based approach [9] to obtain user interest representations. Given the multimodal similarity scores, the score range from 1.0-1.0 to $1.0$ is uniformly partitioned into $N$ predefined intervals. For each interval, the number of similarity scores falling within that range is counted. As a result, an $N$-dimensional histogram vector is constructed, where each dimension corresponds to the frequency of scores in the respective bucket. Then, the $N$-dimensional histogram vector is fed into a subsequent multilayer perceptron for further processing. For modality-enriched modeling, we exploit DTA to obtain user interest representations, detailed in Section IV-C.
@@
The final representation $\mathbf{r}_{u}$ is obtained by aggregating the user interest representations from modality-centric and modality-enriched modeling, respectively, as shown in Eq. (24). While modality-centric modeling strategies emphasize semantic coherence through content-driven generalization, independent of item identities, modality-enriched modeling strategies capture fine-grained behavioral preferences by integrating identity signals with multimodal context, thereby prioritizing context-aware personalization.
Let $\mathbf{r}_{\mathrm{me}}=\mathbf{U}^{\mathrm{DTA}}_{i}\in\mathbb{R}^{d_{h}}$ denote the modality-enriched user interest representation from DTA, where $\mathbf{U}^{\mathrm{DTA}}_{i}$ is the $i$-th row of $\mathbf{U}^{\mathrm{DTA}}$, and let $\mathbf{r}_{\mathrm{mc}}\in\mathbb{R}^{d_{h}}$ denote the modality-centric representation derived from the similarity histogram through an MLP. The final user representation $\mathbf{r}_{u}\in\mathbb{R}^{d_{h}}$ is obtained via a convex combination of the two complementary representations:
$\mathbf{r}_{u}=\alpha\mathbf{r}_{\mathrm{me}}+(1-\alpha)\mathbf{r}_{\mathrm{mc}},$ (24)
where $\alpha\in[0,1]$ controls the trade-off between fine-grained personalization and semantic generalization.
IV-E Discussion
Relationship with KV Caching. DTA is orthogonal to existing KV caching strategies [36] and addresses a distinct challenge in efficient attention computation. Conventional KV caching assumes that key and value representations are target-agnostic, which enables cache and reuse across target items. However, direct integration of target-aware features renders the resulting key and value matrices target-aware, thereby invalidating caching. DTA overcomes this limitation via structural decoupling. It isolates the target-agnostic ID-based pathway, which yields $\mathbf{K}_{\mathrm{id}}$ and $\mathbf{V}_{\mathrm{id}}$, from the target-aware multimodal enhancement. By encoding multimodal similarity scores through discretized embedding lookup rather than instance-specific linear projections, DTA preserves the reusability of base attention components while retaining expressive behavior-semantics interaction within the attention layer. Consequently, DTA does not merely coexist with KV caching. It enables effective caching in modality-enriched sequential recommendation, a regime where standard fusion approaches fail to support computational reuse.
Computational Complexity Comparison of Fusion Strategies. To further quantify the efficiency advantage of DTA, we analyze the floating-point operations (FLOPs) required for computing the key ($\mathbf{K}$) and value ($\mathbf{V}$) matrices, which are the most computationally intensive part of attention [15, 13]. We assume the number of target items is $B$, the user interaction sequence length is $L$, and the embedding dimension before projection is $d_{m}$, with attention hidden size $d_{h}$. For simplicity and fair comparison, we set $d_{m}=d_{h}=d$. As shown in Table I, we compare three fusion strategies: early fusion, late fusion, and our proposed decoupled fusion, under two settings: (1) leveraging the reusability property of target-agnostic nodes, and (2) with such optimization enabled. The FLOPs are derived from the linear projections $\mathbf{W}_{k},\mathbf{W}_{v}\in\mathbb{R}^{d\times d}$, where each projection of an $L\times d$ matrix costs $2Ld^{2}$ FLOPs (accounting for multiply-add operations).
TABLE I: FLOPs of TA leveraging the reusability property of target-agnostic nodes or not for $\mathbf{K}$ and $\mathbf{V}$ computation in different fusion strategies. We assume $B\gg L$ and $d_{m}=d_{h}=d$.
Fusion Strategy Reuse No Reuse
Early Fusion $2BLd^{2}+2BLd$ $2BLd^{2}+2BLd$
Late Fusion $2Ld^{2}+2BLd$ $2BLd^{2}+2BLd$
Decoupled Fusion $2Ld^{2}+2BLd$ $2BLd^{2}+2BLd$
In early fusion, both ID-based embeddings and similarity scores are fused before projection, rendering the resulting $\mathbf{K}$ and $\mathbf{V}$ fully target-aware. Consequently, no component can be reused across target items, and the FLOPs remain identical under both Reuse and No Reuse settings. In contrast, late fusion separates modality processing: the ID-based projections $\mathbf{K}_{\mathrm{id}}$ and $\mathbf{V}_{\mathrm{id}}$ are target-agnostic and thus reusable, reducing the dominant $2BLd^{2}$ term to $2Ld^{2}$ when reuse is enabled.
Our decoupled fusion achieves the same computational efficiency as late fusion under the Reuse setting—since $\mathbf{K}_{\mathrm{id}}$ and $\mathbf{V}_{\mathrm{id}}$ are computed once—while avoiding linear projections on the target-aware similarity scores altogether. Instead, similarity scores are discretized and mapped via embedding lookup, which incurs negligible computational cost (linear in $L$ and independent of $B$) and is omitted in the FLOPs count as it is dominated by projection terms. Therefore, the optimized FLOPs for decoupled fusion match those of late fusion, but with richer cross-modality interaction preserved within the attention mechanism.
To illustrate the practical impact, consider a realistic industrial scenario with $B=1000$ candidate items, $L=400$ historical interactions, and $d=128$. The FLOPs for $\mathbf{K}$$\mathbf{V}$ computation are:
  • Early Fusion: $2\times 1000\times 400\times 128^{2}+2\times 1000\times 400\times 128\approx 13.10$ GFLOPs (identical under both reuse settings).
  • Late Fusion & Decoupled Fusion (with reuse): $2\times 400\times 128^{2}+2\times 1000\times 400\times 128\approx 0.13$ GFLOPs.